# Reducing Energy Bloat in Large Model Training

### v2

@@ C.3. NP-hardness Proof
Appendix D Full Details of GetNextPareto
@@ D.1. Generating Capacity DAG
This equivalence was discovered by Phillips and Dessouky [53, 27].
@@
D.2. Max Flow Algorithm on the Capacity DAG
@@ 35       $c(vu)\leftarrow f(vu)-l(vu)$
Appendix E Proof for Polynomial Runtime
Perseus enumerates the entire “iteration time–energy” Pareto frontier one by one, and the runtime of one iteration is polynomial time with respect to the number of stages $N$ and the number of microbatches $M$. Thus, determining whether the entire algorithm runs in polynomial time reduces to whether the worst case number of iterations is polynomial with respect to $N$ and $M$. While for general DAGs the maximum number of points on the Pareto frontier can be exponential with respect to the size of the DAG [59], here we prove that under mild assumptions for DAGs that represent pipeline schedules, the number of iterations is $O(N+M)$. The assumptions are valid for all pipeline schedules known to the authors, including GPipe [29] and 1F1B [46].
Theorem E.1.
@@ Appendix E Proof for Polynomial Runtime
Appendix F Perseus Client Integration Example
Listing E shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
@@ Appendix F Perseus Client Integration Example
Appendix G Iteration Time–Energy Frontiers
@@
Figure 11 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
@@
Figure 12 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs. Wide-ResNet has a noisy frontier because the variability in microbatch loading time introduces noise in the end-to-end iteration time when computations are tightly packed by Perseus. This was not pronounced in A40 GPUs because compared to A100 PCIe, computation is slower, but data loading time is similar. Thus, the noise in data loading time becomes more noticeable in A100 PCIe.

### v3

@@ C.3. NP-hardness Proof
Appendix D Continuous Relaxation
Figure 11. Pareto-optimal (time, energy) choices for forward and backward computations in each stage of GPT-3 0.3B on NVIDIA A40 GPUs.
The Pipeline Energy Minimization problem is NP-Hard, as we have proved in Appendix C. Thus, we perform continuous relaxation for the problem by fitting an exponential function to each Pareto-optimal (time, energy) points for each forward and backward computation in each stage. Furthermore, the exponential function captures the nature of diminishing returns well, in that the amount of energy consumption needed to reduce computation time by a constant amount increases multiplicatively.
Figure 11 visualizes Pareto-optimal (time, energy) measurements for GPT-3 0.3B on A40 GPUs, showing that the exponential function is a natural fit for data. This was our consistent observation across different GPUs and models in our evaluation.
Appendix E Full Details of GetNextSchedule
@@ E.1. Generating Capacity DAG
This equivalence was given by the theoretical works of Phillips and Dessouky [64, 31].
@@
E.2. Max Flow Algorithm on the Capacity DAG
@@ 35       $c(vu)\leftarrow f(vu)-l(vu)$
Appendix F Proof for Polynomial Runtime
Perseus enumerates the entire time–energy frontier one by one, and the runtime of one iteration is polynomial time with respect to the number of stages $N$ and the number of microbatches $M$. Thus, determining whether the entire algorithm runs in polynomial time reduces to whether the worst case number of iterations is polynomial with respect to $N$ and $M$. While for general DAGs the maximum number of points on the frontier can be exponential with respect to the size of the DAG [71], here we prove that under mild assumptions for DAGs that represent pipeline schedules, the number of iterations is $O(N+M)$. The assumptions are valid for all pipeline schedules known to the authors, including GPipe [34] and 1F1B [56].
Theorem F.1.
@@ Appendix F Proof for Polynomial Runtime
Appendix G Perseus Client Integration Example
Listing F shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
@@ Appendix G Perseus Client Integration Example
Appendix H Time–Energy Frontiers
@@
Figure 12 shows the time–energy frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
@@
Figure 13 shows the time–energy frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs. Wide-ResNet has a noisy frontier because the variability in microbatch loading time introduces noise in the end-to-end iteration time when computations are tightly packed by Perseus. This was not pronounced in A40 GPUs because compared to A100 PCIe, computation is slower, but data loading time is similar. Thus, the noise in data loading time becomes more noticeable in A100 PCIe.